### v1

# High-dimensional components of ${\rm SL}(2,\mathbb{C})$-character varieties of prime knots

Abstract
For a prime knot $K$, we give sufficient conditions for the existence of a component $\mathcal{C}$ of the irreducible ${\rm SL}(2,\mathbb{C})$-character variety of $K$ with $\dim\mathcal{C}>1$, and give a lower bound for $\dim\mathcal{C}$. Specifically, we improve a result of Paoluzzi and Porti on Montesinos knots, and positively answer a question posed by Culler and Dunfield in 2018.
Keywords: ${\rm SL}(2,\mathbb{C})$-character variety; irreducible representation; high-dimensional component; prime knot
MSC2020: 57K10, 57K31
1 Introduction
Let $G={\rm SL}(2,\mathbb{C})$. For a finitely presented group $\Gamma$, call a homomorphism $\rho:\Gamma\to G$ a $G$-representation. Define the character of $\rho$ to be the function $\chi_{\rho}:\Gamma\to\mathbb{C}$, $x\mapsto{\rm tr}(\rho(x))$. Call $\rho$ reducible if the elements of ${\rm Im}(\rho)$ have a common eigenvector; otherwise, call $\rho$ irreducible. Call representations $\rho,\rho^{\prime}$ conjugate and denote $\rho\sim\rho^{\prime}$ if there exists $\mathbf{a}\in G$ such that $\rho^{\prime}(x)=\mathbf{a}\rho(x)\mathbf{a}^{-1}$ for all $x\in\Gamma$. It is known [10] that if $\rho,\rho^{\prime}$ are irreducible, then $\chi_{\rho}=\chi_{\rho^{\prime}}$ if and only if $\rho\sim\rho^{\prime}$. Call $\mathcal{R}(\Gamma):=\hom(\Gamma,G)$ the $G$-representation variety of $\Gamma$. Let $\mathcal{R}^{\rm irr}(\Gamma)=\{\text{irreducible\ }\rho\in\mathcal{R}(\Gamma)\}$. The set $\mathcal{X}(\Gamma)=\{\chi_{\rho}\colon\rho\in\mathcal{R}(\Gamma)\}$ turns out to be an algebraic set, and is called the $G$-character variety of $\Gamma$. The subset $\mathcal{X}^{\rm irr}(\Gamma)=\{\chi_{\rho}\colon\rho\in\mathcal{R}^{\rm irr}(\Gamma)\}$ is Zariski open in $\mathcal{X}(\Gamma)$, and is called the irreducible character variety.
For a $3$-manifold $M$, abbreviate $\mathcal{X}(\pi_{1}(M))$ to $\mathcal{X}(M)$. For a knot $K\subset S^{3}$, let $E_{K}$ denote its exterior, and let $\pi(K)=\pi_{1}(E_{K})$; abbreviate $\mathcal{R}(\pi(K))$ to $\mathcal{R}(K)$ and call it the $G$-representation variety of $K$, and so forth.
Within the character variety of a knot, we have special interests in high-dimensional components (HDCs for short), meaning those of dimension >1>1, for several reasons. Firstly, by the Culler-Shalen theory [10], for a 3-manifold $M$, each ideal point of an algebraic curve in $\mathcal{X}(M)$ is associated to an incompressible surface in $M$. If $\dim\mathcal{X}(K)>1$, then a closed incompressible surface in $E_{K}$ is detected (see [7, Proposition 2.4]). Note that if $M$ is a closed 3-manifold obtained by a Dehn filling of $E_{K}$ for such $K$, then $\dim\mathcal{X}(M)>0$. Secondly, the ${\rm SL}(2,\mathbb{C})$ Casson invariant [8] only uses isolated points in the character variety. It will be amazing if HDCs are taken into account. Thirdly, a conjecture suggested by mathematical physics asserts a surprising identity of adjoint Reidemeister torsions [15], and calls for a better understanding about HDC. Fourthly, HDC plays an important role in the study of Kauffman bracket skein module (KBSM) which has become a central object in quantum topology. By [1], if $\dim\mathcal{X}(M)>0$, then the KBSM of $M$, denoted by $\mathcal{S}(M)$, is not finitely generated. Echoing this, recently it was shown [11] that if $M$ is a closed $3$-manifold with $\dim\mathcal{X}(M)>0$, then $\mathcal{S}(M)$ is not “tame” (in the sense of [11, Theorem 3.1 and Proposition 3.2]). In [5] we found a new type of torsion $e$ in $\mathcal{S}(E_{K})$ for a $4$-strand Montesinos knot $K$; the proof of $e\neq 0$ was built on investigating a HDC of $\mathcal{X}(K)$. As a related but more general result, [2, Corollary 2.2] states that if $M$ is closed with $\dim\mathcal{X}(M)>0$, then $\mathcal{S}(M)$ contains torsion.
However, by now HDC still remains mysterious. Other than [12, 6, 14, 15], few results can be found in the literature.
For almost trivial reasons, the character varieties of connected sums of knots always contain HDCs. In this paper, we focus on primes knots.
We give sufficient conditions for a knot $K$ to admit HDCs, meaning that $\mathcal{X}^{\rm irr}(K)$ contains HDCs, and give lower bounds for the dimensions (see Theorem 3.3). Using elementary techniques, we clarify some significant facts that were never mentioned by other people. Then specifically, we improve a result of Paoluzzi and Porti on the dimensions of character varieties of a class of Montesinos knots (see Remark 3.5).
We work out details for the $(3,3,3,4)$-pretzel knot and two $12$-crossing knots, all of which are hyperbolic. In particular, we positively answer the following question posed by Culler and Dunfield in 2018:
Question 1.1 (Question (7) in [9, Section 9]).
Does there exist a closed atoroidal $3$-manifold $M$ such that $\dim H_{1}(M;\mathbb{Q})\leq 1$ and $\dim\mathcal{X}^{\rm irr}(M)\geq 1$. What if one restricts to 0-surgery on a knot in $S^{3}$?
Based on the experience, we guess that $\dim\mathcal{X}(K)=1$ for any prime knot $K$ with at most $11$ crossings.
2 Preliminary
Most notations adopted in this paper had been introduced in [3, Section 2] and [4, Section 2]. The reader may refer there if necessary.
We always use bold letters to denote elements of $G$. Let $\mathbf{e}$ denote the identity matrix. For $\mathbf{a},\mathbf{b}\in G$, let $\mathbf{a}\lrcorner\mathbf{b}=\mathbf{a}\mathbf{b}\mathbf{a}^{-1}$.
For $t\in\mathbb{C}$, let $\mathbb{B}_{t}=\mathbb{C}\setminus\{2,t^{2}-2\}$, let $G(t)=\{\mathbf{x}\in G\setminus\{\pm\mathbf{e}\}\colon{\rm tr}(\mathbf{x})=t\}.$
Call a tuple $\underline{\mathbf{a}}=(\mathbf{a}_{1},\ldots,\mathbf{a}_{k})\in G^{k}$ reducible if $\mathbf{a}_{1},\ldots,\mathbf{a}_{k}$ share an eigenvector; call $\underline{\mathbf{a}}$ irreducible otherwise. Say that $\underline{\mathbf{a}}^{\prime}=(\mathbf{a}^{\prime}_{1},\ldots,\mathbf{a}^{\prime}_{k})$ is conjugate to $\underline{\mathbf{a}}$ if there exists $\mathbf{c}\in G$ with $\mathbf{c}\lrcorner\underline{\mathbf{a}}=\underline{\mathbf{a}}^{\prime}$, i.e. $\mathbf{c}\lrcorner\mathbf{a}_{i}=\mathbf{a}^{\prime}_{i}$ for $1\leq i\leq k$.
Let
$f_{t}(r_{1},r_{2},r_{3};r)=\$$r^{2}+t(t^{2}-r_{1}-r_{2}-r_{3})r+t^{2}(3-r_{1}-r_{2}-r_{3})$
$+r_{1}^{2}+r_{2}^{2}+r_{3}^{2}+r_{1}r_{2}r_{3}-4.$
The following was stated in [5, Lemma 3.1]:
Lemma 2.1.
(i) $(\mathbf{a}_{1},\mathbf{a}_{2})\in G(t)^{2}$ is irreducible if and only if ${\rm tr}(\mathbf{a}_{1}\mathbf{a}_{2})\in\mathbb{B}_{t}$.
(ii-1) Given $t_{12}\in\mathbb{B}_{t}$, up to conjugacy there exists a unique pair $(\mathbf{a}_{1},\mathbf{a}_{2})\in G(t)^{2}$ with ${\rm tr}(\mathbf{a}_{1}\mathbf{a}_{2})=t_{12}$.
(ii-2) If furthermore $t_{13},t_{23},t_{123}$ are given with $f_{t}(t_{12},t_{13},t_{23};t_{123})=0$, then there exists a unique $\mathbf{a}_{3}\in G(t)$ such that ${\rm tr}(\mathbf{a}_{1}\mathbf{a}_{3})=t_{13}$, ${\rm tr}(\mathbf{a}_{2}\mathbf{a}_{3})=t_{23}$, ${\rm tr}(\mathbf{a}_{1}\mathbf{a}_{2}\mathbf{a}_{3})=t_{123}$.
Remark 2.2.
In (ii-2), we can explicitly determine $\mathbf{a}_{3}$ as follows. Let $s_{0}=t^{2}/2-2$. Let $\overline{\mathbf{a}}_{i}=\mathbf{a}_{i}-(t/2)\mathbf{e}$, let $s_{ij}={\rm tr}(\overline{\mathbf{a}}_{i}\overline{\mathbf{a}}_{j})$, and $s_{123}={\rm tr}(\overline{\mathbf{a}}_{1}\overline{\mathbf{a}}_{2}\overline{\mathbf{a}}_{3})$. Note that $s_{ii}=s_{0}$. Then $f_{t}(t_{12},t_{13},t_{23};t_{123})=0$ is equivalent to
$2s_{123}^{2}+\det((s_{ij})_{3\times 3})=0.$
From ${\rm tr}(\overline{\mathbf{a}}_{1}\overline{\mathbf{a}}_{3})=s_{13}$, ${\rm tr}(\overline{\mathbf{a}}_{2}\overline{\mathbf{a}}_{3})=s_{23}$, ${\rm tr}(\overline{\mathbf{a}}_{1}\overline{\mathbf{a}}_{2}\overline{\mathbf{a}}_{3})=s_{123}$ we can solve
$\overline{\mathbf{a}}_{3}=\frac{1}{s_{0}^{2}-s_{12}^{2}}\big((s_{0}s_{13}-s_{12}s_{23})\overline{\mathbf{a}}_{1}+(s_{0}s_{23}-s_{12}s_{13})\overline{\mathbf{a}}_{2}+2s_{123}\overline{\mathbf{a}}_{1}\overline{\mathbf{a}}_{2}-s_{12}s_{123}\mathbf{e}\big).$
For a knot $K$, in the sprit of Wirtinger presentation, a representation of $K$ identifies with a map $\rho$ assigning to each directed arc $\mathsf{a}$ an element $\rho(\mathsf{a})\in G$ such that $\rho(\mathsf{a}^{-1})=\rho(\mathsf{a})^{-1}$ (where $\mathsf{a}^{-1}$ is obtained by reversing the direction of $\mathsf{a}$), and $\rho(\mathsf{c})=\rho(\mathsf{a})\lrcorner\rho(\mathsf{b})$ for any $\mathsf{a},\mathsf{b},\mathsf{c}$ forming a crossing as
To present such $\rho$, it suffices to choose a direction for each arc and label the arc with an element of $G$. For $t\in\mathbb{C}$, let $\mathcal{R}_{t}(K)$ denote the set of $\rho$ such that $\rho(\mathsf{a})\in G(t)$ for each $\mathsf{a}$, and let $\mathcal{R}_{t}^{\rm irr}(K)$ denote the subset of the irreducible ones; let $\mathcal{X}_{t}^{\rm irr}(K)=\mathcal{R}_{t}^{\rm irr}(K)/\sim$.
In this way, we can talk about representations of tangles. Given a tangle $T$, let $\mathcal{R}(T)$ denote the set of all representations of $T$, and so forth.
Given $\rho\in\mathcal{R}(T)$ and $\mathbf{c}\in G$, let $\mathbf{c}\lrcorner\rho\in\mathcal{R}(T)$ send $\mathsf{a}$ to $\mathbf{c}\lrcorner\rho(\mathsf{a})$. Call $\rho,\rho^{\prime}$ conjugate and denote $\rho\sim\rho^{\prime}$ if $\rho^{\prime}=\mathbf{c}\lrcorner\rho$ for some $\mathbf{c}$.
If $S$ is a subtangle of $T$, then by restriction each $\rho\in\mathcal{R}(T)$ gives rise to a representation of $S$, which is denoted by $\rho|_{S}$.
Figure 1: From left to right: a tangle $T$ in $\mathcal{T}^{2}_{2}$; $N(T)$; $D(T)$; $T_{1}\ast T_{2}$; $T_{1}+T_{2}$.
Let $\mathcal{T}^{2}_{2}$ denote the set of tangles with four ends. For $T\in\mathcal{T}^{2}_{2}$, let $T^{\rm nw}$, $T^{\rm ne}$, $T^{\rm sw}$, $T^{\rm se}$ respectively denote the northwest-, northeast-, southwest- and southeast ends, all directed outward. From $T$ one may construct two links: the numerator $N(T)$ and the denominator $D(T)$. Defined on $\mathcal{T}^{2}_{2}$ are the vertical composition $\ast$ and horizontal composition ++. See Figure 1.
For $p/q\in\mathbb{Q}$, let $[p/q]$ denote the rational tangle associated to $p/q$ (see [3, Section 2.1] for the precise definition). Let $B(p/q)$ denote the rational link $N([p/q])$. Given $p_{1}/q_{1},\ldots,p_{m}/q_{m}\in\mathbb{Q}$, denote $D([p_{1}/q_{1}]\ast\cdots\ast[p_{m}/q_{m}])$ by $M(p_{1}/q_{1},\ldots,p_{m}/q_{m})$ and call it a Montesinos link.
Given $\rho\in\mathcal{R}(T)$ for $T\in\mathcal{T}^{2}_{2}$, let
$\mathbf{g}_{\rho}$ $=\rho(T^{\rm nw})\rho(T^{\rm ne}),\qquad$ ${\rm tr}_{h}(\rho)$ $={\rm tr}(\mathbf{g}_{\rho});$
$\dot{\mathbf{g}}_{\rho}$ $=\rho(T^{\rm sw})\rho(T^{\rm nw}),\qquad$ ${\rm tr}_{v}(\rho)$ $={\rm tr}(\dot{\mathbf{g}}_{\rho}).$
3 High-dimensional components
3.1 A general structure
Lemma 3.1.
Given $\mathbf{a}\in G(t)$, $r\in\mathbb{B}_{t}$, let $\mathcal{C}_{t}^{r}(\mathbf{a})=\{\mathbf{x}\in G(t)\colon{\rm tr}(\mathbf{a}\mathbf{x})=r\}$. Then $\mathcal{C}_{t}^{r}(\mathbf{a})\cong\mathbb{C}^{\times}$ if $t\neq\pm 2$, and $\mathcal{C}_{t}^{r}(\mathbf{a})\cong\mathbb{C}$ if $t\in\{\pm 2\}$.
Proof.
Let
$\mathbf{d}(\kappa)=\left(\begin{array}[]{cc}\kappa&0\\ 0&\kappa^{-1}\end{array}\right),\qquad\mathbf{p}=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right).$
Suppose $\mathbf{b}\in\mathcal{C}_{t}^{r}(\mathbf{a})$, with
$\mathbf{b}=\left(\begin{array}[]{cc}t/2+a&b\\ c&t/2-a\end{array}\right),\qquad bc=\frac{t^{2}}{4}-1-a^{2}.$
When $t\neq\pm 2$, up to conjugacy we may assume $\mathbf{a}=\mathbf{d}(\kappa)$ with $\kappa+\kappa^{-1}=t$. It follows from ${\rm tr}(\mathbf{a}\mathbf{x})=r$ that $(\kappa-\kappa^{-1})a=r-t^{2}/2$, and then
$bc=\frac{t^{2}}{4}-1-\Big(\frac{r-t^{2}/2}{\kappa-\kappa^{-1}}\Big)^{2}=\frac{(r-2)(r-t^{2}+2)}{4-t^{2}}\neq 0.$
When $t=2\varepsilon$ with $\varepsilon\in\{\pm 1\}$, up to conjugacy we may assume $\mathbf{a}=\varepsilon\mathbf{p}$. From ${\rm tr}(\mathbf{a}\mathbf{x})=r$ we see $c=\varepsilon(r-2)\neq 0$, and $b$ is determined by $a$ via
$b=-\frac{a^{2}}{c}=\frac{a^{2}}{\varepsilon(2-r)}.$
Lemma 3.2.
Given $\mathbf{a},\mathbf{b}\in G(t)$ and $\underline{a}=(a_{1},\ldots,a_{n})\in\mathbb{B}_{t}^{n}$ with $n\geq 3$, put
$\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})=\big\{(\mathbf{x}_{1},\ldots,\mathbf{x}_{n-1})\in G(t)^{n-1}\colon{\rm tr}(\mathbf{x}_{i-1}\mathbf{x}_{i})=a_{i},1\leq i\leq n\big\},$
where $\mathbf{x}_{0}=\mathbf{a}$, $\mathbf{x}_{n}=\mathbf{b}$ for convenience. Then $\dim\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})=n-2$.
Proof.
Use induction on $n$ to prove the assertion.
When $n=3$, define
$\Phi:\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})\to\mathcal{C}_{t}^{a_{1}}(\mathbf{a}),\qquad(\mathbf{x}_{1},\mathbf{x}_{2})\mapsto\mathbf{x}_{1}.$
  1. 1.
    If $(\mathbf{x}_{1},\mathbf{b})$ is irreducible, then by Lemma 2.1 and Remark 2.2, there are two or one matrices $\mathbf{x}_{2}\in G(t)$ with ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})=a_{2}$ and ${\rm tr}(\mathbf{x}_{2}\mathbf{b})=a_{3}$.
  2. 2.
    If $(\mathbf{x}_{1},\mathbf{b})$ is reducible and $\mathbf{x}_{1}\mathbf{b}\neq\mathbf{b}\mathbf{x}_{1}$, then $t\neq\pm 2$, and up to conjugacy we may assume
    $\mathbf{x}_{1}=\left(\begin{array}[]{cc}\kappa&0\\ 0&\kappa^{-1}\end{array}\right),\qquad\mathbf{b}=\left(\begin{array}[]{cc}\kappa^{\varepsilon}&1\\ 0&\kappa^{-\varepsilon}\end{array}\right),$
    with $\kappa+\kappa^{-1}=t$, $\varepsilon\in\{\pm 1\}$. For $\mathbf{x}_{2}=(y_{ij})_{2\times 2}\in G(t)$, the conditions ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})=a_{2}$ and ${\rm tr}(\mathbf{b}\mathbf{x}_{2})=a_{3}$ are equivalent to
    $y_{11}=\frac{a_{2}-\kappa^{-1}t}{\kappa-\kappa^{-1}},\qquad y_{22}$ $=\frac{\kappa t-a_{2}}{\kappa-\kappa^{-1}},\qquad y_{21}=a_{3}-\varepsilon a_{2}+\frac{\varepsilon-1}{2}t^{2},$
    $y_{12}y_{21}$ $=y_{11}y_{22}-1=\frac{(a_{2}-2)(a_{2}-t^{2}+2)}{4-t^{2}}.$
    Hence when $a_{3}\neq\varepsilon a_{2}+(1-\varepsilon)t^{2}/2$, there exists a unique $\mathbf{x}_{2}\in G(t)$ with ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})=a_{2}$, ${\rm tr}(\mathbf{b}\mathbf{x}_{2})=a_{3}$; otherwise, no such $\mathbf{x}_{2}$ exists.
  3. 3.
    If $\mathbf{x}_{1}\mathbf{b}=\mathbf{b}\mathbf{x}_{1}$, so that $\mathbf{x}_{1}=\mathbf{b}^{\varepsilon}$ with $\varepsilon\in\{\pm 1\}$, then the space of $\mathbf{x}_{2}$ with ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})=a_{2}$ and ${\rm tr}(\mathbf{b}\mathbf{x}_{2})=a_{3}$ is
    $\begin{cases}\mathcal{C}_{t}^{a_{3}}(\mathbf{b}),&a_{2}=\varepsilon a_{3}+(1-\varepsilon)t^{2}/2\\ \emptyset,&\text{otherwise}\end{cases}.$
By Lemma 3.1, $\dim\mathcal{C}_{t}^{a_{i}}(\mathbf{a})=1$. We have shown $\dim\Phi^{-1}(\mathcal{C}_{t}^{a_{1}}(\mathbf{a})\setminus\{\mathbf{b}^{\pm 1}\})=1$ and $\dim\Phi^{-1}(\mathbf{b}^{\pm 1})\leq 1$. Thus, $\dim\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})=1$.
Suppose $n=k+1\geq 4$ and suppose the assertion holds for $n=k$. For $\underline{a}=(a_{1},\ldots,a_{k+1})$, let $\underline{a}^{\prime}=(a_{1},\ldots,a_{k})$, and define
$\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})\to\mathcal{C}_{t}^{a_{k+1}}(\mathbf{b}),\qquad(\mathbf{x}_{1},\ldots,\mathbf{x}_{k})\mapsto\mathbf{x}_{k}.$
The pre-image of each $\mathbf{x}_{k}\in\mathcal{C}_{t}^{a_{k+1}}(\mathbf{b})$ is $\mathcal{M}_{t}^{\underline{a}^{\prime}}(\mathbf{a},\mathbf{x}_{k})$, which has dimension k2k-2 by the inductive hypothesis. Thus, $\dim\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})=k-1=n-2$. ∎
Theorem 3.3.
Suppose $K=D(T_{0}\ast\cdots\ast T_{n})$, with $T_{0},\ldots,T_{n}\in\mathcal{T}^{2}_{2}$, $n\geq 3$. Suppose for each generic $t$ there exist $\varrho_{i}\in\mathcal{R}_{t}(T_{i})$, $0\leq i\leq n$ such that $\mathbf{g}_{\varrho_{i}}=\mathbf{e}$ for $0\leq i\leq n$, and ${\rm tr}_{v}(\varrho_{i})\neq 2,t^{2}-2$ for $1\leq i\leq n$. Then
$\dim\mathcal{X}^{\rm irr}(K)\geq\begin{cases}n-1,&\varrho_{0}(T_{0}^{\rm ne})\varrho_{0}(T_{0}^{\rm se})\neq\varrho_{0}(T_{0}^{\rm ne})\varrho_{0}(T_{0}^{\rm se})\\ n-2,&\varrho_{0}(T_{0}^{\rm ne})\varrho_{0}(T_{0}^{\rm se})=\varrho_{0}(T_{0}^{\rm ne})\varrho_{0}(T_{0}^{\rm se})\end{cases}.$
Proof.
Let $\mathbf{a}=\varrho_{0}(T_{0}^{\rm se})$, $\mathbf{b}=\varrho_{0}(T_{0}^{\rm ne})^{-1}$. Let $a_{i}=t^{2}-{\rm tr}_{v}(\varrho_{i})$, $1\leq i\leq n$.
Given any $\underline{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n-1})\in\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})$, let $\mathbf{x}_{0}=\mathbf{a}$, $\mathbf{x}_{n}=\mathbf{b}$.
For each $1\leq i\leq n$, since ${\rm tr}_{v}(\varrho_{i})={\rm tr}(\mathbf{x}_{i-1}^{-1}\mathbf{x}_{i})=t^{2}-a_{i}\in\mathbb{B}_{t}$, by Lemma 2.1 (ii-1) there exists $\mathbf{c}_{i}\in G$ such that
$\mathbf{c}_{i}\lrcorner(\varrho_{i}(T_{i}^{\rm ne}),\varrho_{i}(T_{i}^{\rm se}))=(\mathbf{x}_{i-1}^{-1},\mathbf{x}_{i}).$
Put $\rho_{i}=\mathbf{c}_{i}\lrcorner\varrho_{i}\in\mathcal{R}_{t}(T_{i})$. Then $\rho_{i}(T_{i}^{\rm ne})=\mathbf{x}_{i-1}^{-1}$, $\rho_{i}(T_{i}^{\rm se})=\mathbf{x}_{i}$, and $\mathbf{g}_{\rho_{i}}=\mathbf{e}$.
Let $\rho_{0}=\varrho_{0}$. Due to the compatibilities of the values at common arcs,
$\rho_{i-1}(T_{i-1}^{\rm se})=\rho_{i}(T_{i}^{\rm ne})^{-1}\ \ \ (1\leq i\leq n),\qquad\rho_{n}(T_{n}^{\rm se})=\rho_{0}(T_{0}^{\rm ne})^{-1},$
we can glue $\rho_{0},\ldots,\rho_{n}$ into a representation $\rho^{\underline{\mathbf{x}}}\in\mathcal{R}_{t}^{\rm irr}(K)$.
This defines a map $\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})\to\mathcal{R}_{t}^{\rm irr}(K)$.
Suppose $\mathbf{a}\mathbf{b}\neq\mathbf{b}\mathbf{a}$. If $\rho^{\underline{\mathbf{x}}}\sim\rho^{\underline{\mathbf{x}}^{\prime}}$, say $\rho^{\underline{\mathbf{x}}^{\prime}}=\mathbf{c}\lrcorner\rho^{\underline{\mathbf{x}}}$ for some $\mathbf{c}\in G$, then $\mathbf{c}\lrcorner(\mathbf{a},\mathbf{b})=(\mathbf{a},\mathbf{b})$, so $\mathbf{c}=\pm\mathbf{e}$, implying $\underline{\mathbf{x}}=\underline{\mathbf{x}}^{\prime}$. Hence $\dim\mathcal{X}_{t}^{\rm irr}(K)\geq n-2$ for each generic $t$. Thus, $\dim\mathcal{X}^{\rm irr}(K)\geq n-1$.
Now suppose $\mathbf{a}\mathbf{b}=\mathbf{b}\mathbf{a}$, so that $\mathbf{b}\in\{\mathbf{a}^{\pm 1}\}$. Fix $\mathbf{v}\in\mathcal{C}_{t}^{a_{1}}(\mathbf{a})$. For any $\underline{\mathbf{x}}^{\prime}=(\mathbf{x}^{\prime}_{1},\ldots,\mathbf{x}^{\prime}_{n-1})\in\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})$, since ${\rm tr}(\mathbf{a}\mathbf{x}^{\prime}_{1})={\rm tr}(\mathbf{a}\mathbf{v})\in\mathbb{B}_{t}$, by Lemma 2.1 (ii-1) we can find $\mathbf{c}\in G$ such that $\mathbf{c}\lrcorner(\mathbf{a},\mathbf{x}^{\prime}_{1})=(\mathbf{a},\mathbf{v})$. Note that $\mathbf{c}\lrcorner\mathbf{b}=\mathbf{b}$. Hence $\rho^{\underline{\mathbf{x}}^{\prime}}\sim\rho^{\underline{\mathbf{x}}}$ for some $\underline{\mathbf{x}}$ with $\mathbf{x}_{1}=\mathbf{v}$. This shows $\dim\mathcal{X}_{t}^{\rm irr}(K)\geq n-3$ for each generic $t$. Thus, $\dim\mathcal{X}^{\rm irr}(K)\geq n-2$. ∎
Corollary 3.4.
Let $K=M(\beta_{0}/\alpha_{0},\ldots,\beta_{n}/\alpha_{n})$ be a Montesinos knot with $n\geq 3$. If each $\alpha_{i}\geq 3$, then $\dim\mathcal{X}^{\rm irr}(K)\geq n-1$.
Proof.
According to [13, Section 7] (also see [5, Section 3.1]), for each generic $t$, there exists $\sigma_{i}\in\mathcal{R}^{\rm irr}(B(\beta_{i}/\alpha_{i}))$, which identifies with some $\varrho_{i}\in\mathcal{R}_{t}(T_{i})$ such that $\mathbf{g}_{\varrho_{i}}=\mathbf{e}$ and ${\rm tr}_{v}(\varrho_{i})\neq 2,t^{2}-2$. In particular, $\varrho_{0}(T_{0}^{\rm ne})\varrho_{0}(T_{0}^{\rm se})\neq\varrho_{0}(T_{0}^{\rm ne})\varrho_{0}(T_{0}^{\rm se})$. The assertion follows from Theorem 3.3. ∎
Remark 3.5.
On restriction to the case $2\mid\alpha_{0}$ (and $2\nmid\alpha_{i}$ for $1\leq i\leq n$), this improves the result of [14, Theorem 1] that $\dim\mathcal{X}(K)\geq n-2$.
For a connected sum of knots, essentially we achieve the same effect as the “bending trick” does. The key difference is that [14, Section 4] requires $\rho(\mu_{1})\rho(\mu_{n+1})=\rho(\mu_{n+1})\rho(\mu_{1})$. The commuting condition decreases the dimension by $1$, as seen from the last paragraph of the proof of $3.$
Figure 2: Left: the knot $K_{n}$. Right: $N(T_{0})=B(-(4n+1)/n)$.
Example 3.6 (A non-Montesinos knot).
Let $K_{n}=D(T_{0}\ast T_{1}\ast T_{2}\ast T_{3})$, with $T_{0}=([n]\ast[1/2])+[1/2]$, and $T_{i}=[p_{i}/q_{i}]$, $i=1,2,3$. See Figure 2, left. Suppose $q_{1},q_{2},q_{3}\geq 3$, and $n\geq 3$.
Note that $N(T_{0})=B(-(4n+1)/n)$, as illustrated in Figure 2, right. Given each generic $t\neq 0$, there exists $\rho_{0}\in\mathcal{R}_{t}^{\rm irr}(B(-\frac{4n+1}{n}))$, which is the same as some $\varrho_{0}\in\mathcal{R}_{t}(T_{0})$ with $\mathbf{g}_{\varrho_{0}}=\mathbf{e}$. Then $\mathbf{y}:=\varrho_{0}(y)=\varrho_{0}(T_{0}^{\rm ne})$ does not commute with $\mathbf{x}:=\varrho_{0}(x)=\varrho_{0}(T_{0}^{\rm se})$. This is because $\mathbf{y}=\mathbf{z}\mathbf{x}\mathbf{z}^{-1}$, where $\mathbf{z}=\varrho_{0}(z)$; if $\mathbf{x}\mathbf{y}=\mathbf{y}\mathbf{x}$, then $\mathbf{z}\mathbf{x}\mathbf{z}^{-1}\in\{\mathbf{x}^{\pm 1}\}$, so that $\mathbf{z}^{2}\mathbf{x}\mathbf{z}^{-2}=\mathbf{z}\mathbf{y}\mathbf{z}^{-1}=\mathbf{x}$, implying $\mathbf{y}\mathbf{z}=\mathbf{z}\mathbf{y}$, which contradicts the irreducibility of $\rho_{0}$.
For each $1\leq i\leq 3$, similarly as in the proof of Corollary 3.4, there exists $\varrho_{i}\in\mathcal{R}_{t}(T_{i})$ such that $\mathbf{g}_{\varrho_{i}}=\mathbf{e}$ and ${\rm tr}_{v}(\varrho_{i})\neq 2,t^{2}-2$.
By Theorem 3.3, $\dim\mathcal{X}^{\rm irr}(K_{n})\geq 2$.
A nontrivial answer to the following will be very interesting:
Question 3.7.
Is there a prime knot $K$ not of the form $D(T_{0}\ast\cdots\ast T_{n})$ with $T_{i}\in\mathcal{T}_{2}^{2}$ and $n\geq 3$ that admits a HDC?
3.2 Conditional structures
We partially answer Question 3.7, by showing that when
$K=D(T_{1}\ast T_{2}\ast(T_{3}+T_{4}))$
for some $T_{i}\in\mathcal{T}_{2}^{2}$, there are two possibilities for $K$ to admit a HDC. In each case, the $T_{i}$’s are required to satisfy some conditions.
Figure 3:
  1. Case 1
    Suppose that for each generic $t$, there exist $\varrho_{i}\in\mathcal{R}_{t}(N(T_{i}))$, $i=1,2$ and $\varrho_{j}\in\mathcal{R}_{t}(D(T_{j}))$, $j=3,4$ such that ${\rm tr}_{v}(\varrho_{1})={\rm tr}_{v}(\varrho_{2})=a\in\mathbb{B}_{t}$, and ${\rm tr}_{h}(\varrho_{3})={\rm tr}_{h}(\varrho_{4})=b\in\mathbb{B}_{t}$. This requires $T_{1},T_{2}$ to be related in a certain way, and similarly for $T_{3},T_{4}$.
    Given $(\mathbf{x},\mathbf{y},\mathbf{z})\in G(t)^{3}$ with ${\rm tr}(\mathbf{x}\mathbf{y})=a$, ${\rm tr}(\mathbf{x}\mathbf{z})=b$, there exists $\rho\in\mathcal{R}_{t}(K)$ with $\rho(x_{1})=\rho(x_{2})=\rho(x_{3})=\rho(x_{4})=\mathbf{x}$, $\rho(y_{1})=\rho(y_{2})=\mathbf{y}$, $\rho(z_{1})=\rho(z_{2})=\mathbf{z}$. See the left part of Figure 3 for demonstration.
    The conjugacy classes of $(\mathbf{x},\mathbf{y},\mathbf{z})$ are parameterized by $c:={\rm tr}(\mathbf{y}\mathbf{z})$ and $r:={\rm tr}(\mathbf{x}\mathbf{y}\mathbf{z})$, subject to $f_{t}(a,b,c;r)=0$. Therefore, $K$ admits a HDC.
  2. Case 2
    Suppose $T_{4}=\sigma(T_{3})$, where $\sigma$ is the reflection along the dotted line (see the right part of Figure 3), so $\mathsf{a}\mapsto\mathsf{a}^{\prime}:=\sigma(\mathsf{a})^{-1}$ establishes a bijection between directed arcs of $T_{3}$ and those of $T_{4}$. Furthermore, suppose that for each generic $t$, there exists $\varrho_{i}\in\mathcal{R}_{t}(N(T_{i}))$ with ${\rm tr}_{v}(\varrho_{i})=a_{i}\in\mathbb{B}_{t}$, $i=1,2$, and for generic $b$, there exists $\tau\in\mathcal{R}_{t}(T_{3})$ with ${\rm tr}_{v}(\varrho_{i})=b$.
    Given $(\mathbf{x},\mathbf{y},\mathbf{z})\in G(t)^{3}$ with ${\rm tr}(\mathbf{x}\mathbf{y})=a_{1}$, ${\rm tr}(\mathbf{y}\mathbf{z})=a_{2}$, and ${\rm tr}(\mathbf{x}\mathbf{z})$ being generic, by the assumption, there exists $\varrho\in\mathcal{R}(T_{3})$ with $\varrho(x_{1})=\mathbf{x}$, $\varrho(z_{1})=\mathbf{z}$. Put $\varrho^{\prime}(\mathsf{a}^{\prime})=\varrho(\mathsf{a})$ for each $\mathsf{a}^{\prime}$. Then $\varrho^{\prime}(\mathsf{c}^{\prime})=\varrho^{\prime}(\mathsf{a}^{\prime})\lrcorner\varrho^{\prime}(\mathsf{b}^{\prime})$ for any directed arcs $\mathsf{a}^{\prime},\mathsf{b}^{\prime},\mathsf{c}^{\prime}$ making a crossing of $T_{4}$ (see Figure 4), as $\varrho(\mathsf{c})=\varrho(\mathsf{a})\lrcorner\varrho(\mathsf{b})$. So $\varrho^{\prime}\in\mathcal{R}(T_{4})$, which satisfies $\varrho^{\prime}(x_{2})=\mathbf{x}$, $\varrho^{\prime}(z_{2})=\mathbf{z}$. Gluing $\varrho$ and $\varrho^{\prime}$ yields $\rho_{\mathbf{x}}^{\mathbf{z}}\in\mathcal{R}(T_{3}+T_{4})$ with $\rho_{\mathbf{x}}^{\mathbf{z}}(x_{i})=\mathbf{x}$ and $\rho_{\mathbf{x}}^{\mathbf{z}}(z_{i})=\mathbf{z}$ for $i=1,2$. By the assumption again, it can be extended to $\rho\in\mathcal{R}_{t}(K)$ with $\rho(x_{i})=\mathbf{x}$, $\rho(y_{i})=\mathbf{y}$, $\rho(z_{i})=\mathbf{z}$, $i=1,2$.
    The conjugacy classes of $(\mathbf{x},\mathbf{y},\mathbf{z})$ are parameterized by $b:={\rm tr}(\mathbf{x}\mathbf{z})$ and $r:={\rm tr}(\mathbf{x}\mathbf{y}\mathbf{z})$, subject to $f_{t}(a_{1},a_{2},b;r)=0$. Thus, there is a surjection $\mathcal{X}_{t}^{\rm irr}(K)\to\{(b,r)\colon f_{t}(a_{1},a_{2},b;r)=0\}$. Therefore, $K$ admits a HDC.
Figure 4: Left: a crossing in $T_{3}$. Right: the corresponding one in $T_{4}$ under $\sigma$.
4 Examples
In this section, we study three hyperbolic knots, explicitly describing the HDCs they admit. As a byproduct, we positively answer Question 1.1.
For $r=\mu+\mu^{-1}$, let
$\gamma_{k}(r)=\begin{cases}(\mu^{k}-\mu^{-k})/(\mu-\mu^{-1}),&\mu\neq\pm 1\\ k\mu^{k-1},&\mu=\pm 1\end{cases}.$
By Cayley-Hamilton Theorem, if $\mathbf{x}\in G(r)$, then for each $k\in\mathbb{Z}$,
$\mathbf{x}^{k}=\gamma_{k}(r)\mathbf{x}-\gamma_{k-1}(r)\mathbf{e}.$
In particular, if $\mathbf{x}\in G(0)$, then $\mathbf{x}^{2}=-\mathbf{e}$; if $\mathbf{x}\in G(1)$, then $\mathbf{x}^{3}=-\mathbf{e}$.
Lemma 4.1.
If ${\rm tr}(\mathbf{a})={\rm tr}(\mathbf{b})$ and ${\rm tr}(\mathbf{a}\mathbf{b})=1$, then $\mathbf{a}\mathbf{b}\mathbf{a}^{k}\mathbf{b}\mathbf{a}=-\mathbf{b}^{k-2}$ for each $k\in\mathbb{Z}$; in particular, $\mathbf{a}\mathbf{b}^{2}\mathbf{a}=-\mathbf{b}^{-2}$.
Proof.
Since ${\rm tr}(\mathbf{a}\mathbf{b}^{2})={\rm tr}(\mathbf{a}(t\mathbf{b}-\mathbf{e}))=t\cdot{\rm tr}(\mathbf{a}\mathbf{b})-{\rm tr}(\mathbf{a})=0$, we have $(\mathbf{a}\mathbf{b}^{2})^{2}=-\mathbf{e}$, so that $\mathbf{a}\mathbf{b}^{2}\mathbf{a}=-\mathbf{b}^{-2}$.
For the general case, note that $(\mathbf{a}\mathbf{b})^{3}=-\mathbf{e}$, and then
$\mathbf{a}\mathbf{b}\mathbf{a}^{k}\mathbf{b}\mathbf{a}\mathbf{b}^{2}$ $=\mathbf{a}\mathbf{b}(\gamma_{k}(t)\mathbf{a}-\gamma_{k-1}(t)\mathbf{e})\mathbf{b}\mathbf{a}\mathbf{b}^{2}$
$=\gamma_{k}(t)(\mathbf{a}\mathbf{b})^{3}\cdot\mathbf{b}-\gamma_{k-1}(t)\mathbf{a}\mathbf{b}^{2}\mathbf{a}\cdot\mathbf{b}^{2}$
$=-\gamma_{k}(t)\mathbf{b}+\gamma_{k-1}(t)\mathbf{e}=-\mathbf{b}^{k}.$
implying $\mathbf{a}\mathbf{b}\mathbf{a}^{k}\mathbf{b}\mathbf{a}=-\mathbf{b}^{k-2}$. ∎
4.1 The $(3,3,3,4)$-pretzel knot
Let $P$ denote the $(3,3,3,4)$-pretzel knot, as shown in Figure 5, left.
Each quadruple $\underline{\mathbf{x}}=(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\mathbf{x}_{4})\in G(t)^{4}$ satisfying
${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})={\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})={\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})=1,\qquad{\rm tr}(\mathbf{x}_{4}\mathbf{x}_{1})=0$ (1)
determines a representation $\rho_{\underline{\mathbf{x}}}\in\mathcal{R}_{t}^{\rm irr}(K)$ such that $\mathbf{x}_{i}=\rho_{\underline{\mathbf{x}}}(x_{i})$, $1\leq i\leq 4$ and $\mathbf{g}_{\rho_{\underline{\mathbf{x}}}}=\mathbf{e}$.
Figure 5: The $(3,3,3,4)$-pretzel knot.
Claim 4.2.
Let $\mathcal{Y}_{t}$ denote the set of conjugacy classes of $\underline{\mathbf{x}}\in G(t)^{4}$ satisfying (1). The map $\Psi_{t}:\mathcal{Y}_{t}\to\mathbb{C}^{3}$ defined by
$\underline{\mathbf{x}}\mapsto({\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3}),(\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3}),{\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3}\mathbf{x}_{4}))$
is injective, and its image consists of $(t_{13},t_{123},t_{134})$ with $f_{t}(1,1,t_{13};t_{123})=f_{t}(1,0,t_{13};t_{134})=0$, $t_{13}\neq t^{2}-2$, and $t_{13}\neq 2$ if $t^{2}\in\{2,3\}$.
Proof.
Given any $t_{13},t_{123}$ with $f_{t}(1,1,t_{13};t_{123})=0$ and $t_{13}\in\mathbb{B}_{t}$, up to conjugacy there uniquely exists $(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3})\in G(t)^{3}$ such that ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})={\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})=1$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3})=t_{13}$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3})=t_{123}$. Since $t_{13}\neq 2,t^{2}-2$, by Lemma 2.1 (ii-2), any $t_{134}$ with $f_{t}(1,0,t_{13};t_{134})=0$ further determines a unique $\mathbf{x}_{4}$ such that ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})=0$, ${\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})=1$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3}\mathbf{x}_{4})=t_{134}$.
Now assume $t_{13}\in\{2,t^{2}-2\}$. Given any $t_{134}$ with $f_{t}(1,0,t_{13};t_{134})=0$, up to conjugacy there is a unique $(\mathbf{x}_{1},\mathbf{x}_{3},\mathbf{x}_{4})\in G(t)^{3}$ such that ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})=0$, ${\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})=1$, ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{3}\mathbf{x}_{4})=t_{134}$. Since ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})\neq{\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})$, we have $\mathbf{x}_{1}\neq\mathbf{x}_{3}$. Note that $\mathbf{x}_{1}=\mathbf{x}_{3}^{-1}$ if and only if $t^{2}-2={\rm tr}(\mathbf{x}_{1}\mathbf{x}_{4})+{\rm tr}(\mathbf{x}_{3}\mathbf{x}_{4})=1$, i.e. $t^{2}=3$; when this holds, there does not exist $\mathbf{x}_{2}$ with ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})={\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})=1$, as ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})+{\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})\neq t^{2}-2$. Conversely, suppose $(t^{2},t_{13})\neq(3,2)$, then $(\mathbf{x}_{1},\mathbf{x}_{3})$ is reducible and $\mathbf{x}_{1}\mathbf{x}_{3}\neq\mathbf{x}_{3}\mathbf{x}_{1}$. Following the proof of Lemma 3.2, in the context of Point 2 there, we can deduce $\varepsilon=-1$, so that $t_{13}=2$, and $t^{2}\neq 2$. When these hold, there exists $\mathbf{x}_{2}$ satisfying ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})={\rm tr}(\mathbf{x}_{2}\mathbf{x}_{3})=1$; it is fixed by $t_{123}={\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2}\mathbf{x}_{3})$ which satisfies $f_{t}(1,1,2;t_{123})=0$. ∎
Consequently, $P$ admits a HDC isomorphic to
$\big\{(t,t_{13},t_{123},t_{134})\colon$ $t_{13}\neq t^{2}-2\ \text{and\ }(t^{2},t_{13})\neq(2,2),(3,2),$
$t_{123}^{2}+t(t^{2}-2-t_{13})t_{123}+t_{13}^{2}+(1-t^{2})t_{13}+t^{2}-2=0,$
$t_{134}^{2}+t(t^{2}-1-t_{13})t_{134}+t_{13}^{2}-t^{2}t_{13}+2t^{2}-3=0\big\}.$
In general, given an oriented knot $K$, each directed arc provides a meridional generator for $\pi(K)$. Choose a preferred meridian $\mathfrak{m}$. Then the element represented by the longitude $\mathfrak{l}$ paired with $\mathfrak{m}$ can be computed as follows. Starting at the point marked with a bullet, walk along the diagram guided by the orientation of $K$; denote the generator contributed by the over-crossing arc at the $k$-th crossing (labeled by $k$) by $\mathfrak{z}_{k}$. Then $\mathfrak{l}=\mathfrak{m}^{w(K)}\mathfrak{z}_{1}\cdots\mathfrak{z}_{n}$, where $n$ is the number of crossings, and $w(K)$ is the writhe whose presence ensures $\mathfrak{l}$ to represent 0 in $H_{1}(E_{K})$.
For the present knot $P$, choose $x_{1}$ as $\mathfrak{m}$. Note that $w(P)=13$.
For $\underline{\mathbf{x}}\in\mathcal{Y}_{t}$, let $\mathbf{l}=\rho_{\underline{\mathbf{x}}}(\mathfrak{l})$. As illustrated in the right part of Figure 5, $\rho_{\underline{\mathbf{x}}}(\mathfrak{z}_{1})=\mathbf{x}_{4}^{-1}$, $\rho_{\underline{\mathbf{x}}}(\mathfrak{z}_{2})=\mathbf{x}_{1}^{-1}\mathbf{x}_{4}^{-1}\mathbf{x}_{1}$, $\rho_{\underline{\mathbf{x}}}(\mathfrak{z}_{3})=\mathbf{x}_{2}^{-1}$, etc., hence
$\mathbf{l}$ $=\mathbf{x}_{1}^{13}\cdot\mathbf{x}_{4}^{-1}\cdot\mathbf{x}_{1}^{-1}\mathbf{x}_{4}^{-1}\mathbf{x}_{1}\cdot\mathbf{x}_{2}^{-1}\cdot\mathbf{x}_{1}^{-1}\cdot\mathbf{x}_{2}\mathbf{x}_{3}^{-1}\mathbf{x}_{2}^{-1}\cdot\mathbf{x}_{4}^{-1}\cdot\mathbf{x}_{3}^{-1}$
$\ \ \ \ \cdot\mathbf{x}_{4}\mathbf{x}_{1}^{-1}\mathbf{x}_{4}^{-1}\cdot\mathbf{x}_{1}^{-1}\cdot\mathbf{x}_{4}\mathbf{x}_{3}^{-1}\mathbf{x}_{4}^{-1}\cdot\mathbf{x}_{2}^{-1}\cdot\mathbf{x}_{3}^{-1}\cdot\mathbf{x}_{2}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}^{-1}$
$=\mathbf{x}_{1}^{14}(\mathbf{x}_{4}\mathbf{x}_{1})^{-2}\mathbf{x}_{1}\mathbf{x}_{2}^{-1}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}\mathbf{x}_{3}^{-1}\mathbf{x}_{2}^{-1}\mathbf{x}_{4}^{-1}\mathbf{x}_{3}^{-1}\mathbf{x}_{4}(\mathbf{x}_{4}\mathbf{x}_{1})^{-2}\mathbf{x}_{4}^{2}\mathbf{x}_{3}^{-1}\mathbf{x}_{4}^{-1}\mathbf{x}_{2}^{-1}\mathbf{x}_{3}^{-1}\mathbf{x}_{2}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}^{-1}$
$=\mathbf{x}_{1}^{15}\mathbf{x}_{2}^{-1}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}\mathbf{x}_{3}^{-1}\mathbf{x}_{2}^{-1}(\mathbf{x}_{4}^{-1}\mathbf{x}_{3}^{-1}\mathbf{x}_{4}^{3}\mathbf{x}_{3}^{-1}\mathbf{x}_{4}^{-1})\mathbf{x}_{2}^{-1}\mathbf{x}_{3}^{-1}\mathbf{x}_{2}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}^{-1}$
$=-\mathbf{x}_{1}^{15}\mathbf{x}_{2}^{-1}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}(\mathbf{x}_{3}^{-1}\mathbf{x}_{2}^{-1}\mathbf{x}_{3}^{5}\mathbf{x}_{2}^{-1}\mathbf{x}_{3}^{-1})\mathbf{x}_{2}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}^{-1}$
$=\mathbf{x}_{1}^{15}(\mathbf{x}_{2}^{-1}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}^{9}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}^{-1})$
$=-\mathbf{x}_{1}^{26}.$
In the third line, $(\mathbf{x}_{4}\mathbf{x}_{1})^{2}=-\mathbf{e}$ is used, and in the last three lines, Lemma 4.1 is applied.
Let $P(a/b)$ denote the Dehn filling of $E_{P}$ with slope $a/b$. For $\rho$ to descend to $\pi_{1}(P(a/b))\to G$, a sufficient and necessary condition is
$\mathbf{e}=\mathbf{x}_{1}^{a}\mathbf{l}^{b}=(-1)^{b}\mathbf{x}_{1}^{a+26b},$
which is equivalent to $t=\kappa+\kappa^{-1}$, with $\kappa^{a+26b}=(-1)^{b}$ and $\kappa\neq\pm 1$.
By [16, Theorem 3.6], $P(a/b)$ is hyperbolic if $a/b\neq 1/0$. Thus, as long as $a+26b\neq\pm 1$, there exists $t\neq\pm 2$ such that $\mathcal{X}^{\rm irr}(P(a/b))$ contains a component contributed by $\mathcal{X}^{\rm irr}_{t}(P)$ which is positive-dimensional, so $P(a/b)$ provides a positive answer to Question 1.1. In particular, $P(0/1)$ does.
4.2 Two knots with $12$ crossings
Now we consider two more knots $Q_{1},Q_{2}$, to illustrate the two conditional structures for HDCs given in Section 3.2.
Each knot also provides positive answers to Question 1.1, for similar reasons as $P$, which we no longer explain.
Figure 6: The knot $Q_{1}$.
Let $Q_{1}=D([3]\ast[3]\ast([-1/3]+[-1/3]))$, as shown in Figure 6, left.
Each triple $\underline{\mathbf{w}}=(\mathbf{x},\mathbf{y},\mathbf{z})\in G(t)^{3}$ satisfying
${\rm tr}(\mathbf{x}\mathbf{y})={\rm tr}(\mathbf{x}\mathbf{z})=1,$ (2)
determines a representation $\rho_{\underline{\mathbf{w}}}\in\mathcal{R}_{t}^{\rm irr}(Q_{1})$ such that $\rho_{\underline{\mathbf{w}}}(x_{i})=\mathbf{x}$, $1\leq i\leq 4$; $\rho_{\underline{\mathbf{w}}}(y_{1})=\rho_{\underline{\mathbf{w}}}(y_{2})=\mathbf{y}$, $\rho_{\underline{\mathbf{w}}}(z_{1})=\rho_{\underline{\mathbf{w}}}(z_{2})=\mathbf{z}.$
The conjugacy classes of $\underline{\mathbf{w}}\in G(t)^{3}$ with (2) are parameterized by $t_{23}={\rm tr}(\mathbf{y}\mathbf{z})$ and $t_{123}={\rm tr}(\mathbf{x}\mathbf{y}\mathbf{z})$, subject to $f_{t}(1,1,t_{23};t_{123})=0$. Hence $Q_{1}$ admits a HDC isomorphic to
$\big\{(t,t_{23},t_{123})\colon t_{123}^{2}+t(t^{2}-2-t_{23})t_{123}+t_{23}^{2}+(1-t^{2})t_{23}+t^{2}-2=0\big\}.$
Choose $x_{1}$ as the preferred meridian $\mathfrak{m}$. Note that $w(Q_{1})=12$.
Given $\underline{\mathbf{w}}$ with (2), as illustrated in the right part of Figure 6, we compute
$\rho_{\underline{\mathbf{w}}}(\mathfrak{l})$ $=\mathbf{x}^{12}\cdot\mathbf{x}\mathbf{z}^{-1}\mathbf{x}^{-1}\cdot\mathbf{x}^{-1}\cdot\mathbf{z}^{-1}\cdot\mathbf{y}^{-1}\cdot\mathbf{x}^{-1}\cdot\mathbf{x}^{-1}\mathbf{y}^{-1}\mathbf{x}$
$\ \ \ \ \cdot\mathbf{x}\mathbf{z}^{-1}\mathbf{x}^{-1}\cdot\mathbf{x}^{-1}\cdot\mathbf{z}^{-1}\cdot\mathbf{y}^{-1}\cdot\mathbf{x}^{-1}\cdot\mathbf{x}^{-1}\mathbf{y}^{-1}\mathbf{x}$
$=\mathbf{x}^{13}(\mathbf{z}^{-1}\mathbf{x}^{-2}\mathbf{z}^{-1})(\mathbf{y}^{-1}\mathbf{x}^{-2}\mathbf{y}^{-1})\mathbf{x}^{2}(\mathbf{z}^{-1}\mathbf{x}^{-2}\mathbf{z}^{-1})(\mathbf{y}^{-1}\mathbf{x}^{-2}\mathbf{y}^{-1})\mathbf{x}$
$=\mathbf{x}^{13}(-\mathbf{x}^{2})(-\mathbf{x}^{2})\mathbf{x}^{2}(-\mathbf{x}^{2})(-\mathbf{x}^{2})\mathbf{x}$
$=\mathbf{x}^{24}.$
Figure 7: The knot $Q_{2}$.
Let $Q_{2}=D([3]\ast[3]\ast([-1/3]+[1/3]))$, as shown in Figure 7, left.
Suppose $(\mathbf{x},\mathbf{y},\mathbf{z})\in G(t)^{3}$ satisfies
${\rm tr}(\mathbf{x}\mathbf{y})={\rm tr}(\mathbf{y}\mathbf{z})=1.$ (3)
By [4, Lemma 3.2 and Example 3.5], given $r\in\mathbb{B}_{t}$ with
$2-(r+2-t^{2})(r-1)^{2}={\rm tr}(\mathbf{z}^{-1}\mathbf{x})=t^{2}-{\rm tr}(\mathbf{x}\mathbf{z}),$
there exists a unique $\mathbf{v}\in G(t)$ with
${\rm tr}(\mathbf{z}\mathbf{v})=r,\qquad{\rm tr}(\mathbf{x}^{-1}\mathbf{v})=2+(r-2)(r+2-t^{2}).$ (4)
To avoid subtleties, we assume $r\in\mathbb{B}_{t}$. Each quadruple $\underline{\mathbf{w}}=(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{v})$ satisfying (3), (4) determines a representation $\rho_{\underline{\mathbf{w}}}\in\mathcal{R}_{t}^{\rm irr}(Q_{2})$ with $\rho_{\underline{\mathbf{w}}}(x_{i})=\mathbf{x}$, $\rho_{\underline{\mathbf{w}}}(y_{i})=\mathbf{y}$, $\rho_{\underline{\mathbf{w}}}(z_{i})=\mathbf{z}$ for $i=1,2$ and $\rho_{\underline{\mathbf{w}}}(v)=\mathbf{v}$.
The conjugacy classes of $\underline{\mathbf{w}}\in G(t)^{4}$ with (3), (4) are parameterized by $t_{13}={\rm tr}(\mathbf{x}\mathbf{z})$ and $t_{123}={\rm tr}(\mathbf{x}\mathbf{y}\mathbf{z})$, subject to $f_{t}(1,1,t_{13};t_{123})=0$. Hence $Q_{2}$ admits a HDC, a Zariski open subset of which is isomorphic to
$\big\{(t,t_{13},t_{123},r)\colon$ $t_{123}^{2}+t(t^{2}-2-t_{13})t_{123}+t_{13}^{2}+(1-t^{2})t_{13}+t^{2}-2=0,$
$(r+2-t^{2})(r-1)^{2}=t_{13}-t^{2}+2,\ r\neq 2,t^{2}-2\big\}.$
Choose $x_{1}$ as $\mathfrak{m}$. Note that $w(Q_{2})=6$. For $\underline{\mathbf{w}}\in G(t)^{4}$ satisfying (3), (4), as illustrated in the right part of Figure 7, we compute
$\rho_{\underline{\mathbf{w}}}(\mathfrak{l})$ $=\mathbf{x}^{6}\cdot\mathbf{x}\mathbf{u}^{-1}\mathbf{x}^{-1}\cdot\mathbf{x}\mathbf{u}\mathbf{x}^{-1}\cdot\mathbf{y}^{-1}\cdot\mathbf{x}^{-1}\cdot\mathbf{y}\mathbf{z}^{-1}\mathbf{y}^{-1}\cdot\mathbf{v}$
$\ \ \ \ \cdot\mathbf{x}\cdot\mathbf{x}^{-1}\cdot\mathbf{v}^{-1}\cdot\mathbf{y}^{-1}\cdot\mathbf{z}^{-1}\cdot\mathbf{y}\mathbf{x}^{-1}\mathbf{y}^{-1}$
$=\mathbf{x}^{6}\mathbf{y}^{-1}\mathbf{x}^{-1}\mathbf{y}(\mathbf{z}^{-1}\mathbf{y}^{-2}\mathbf{z}^{-1})\mathbf{y}\mathbf{x}^{-1}\mathbf{y}^{-1}$
$=-\mathbf{x}^{6}(\mathbf{y}^{-1}\mathbf{x}^{-1}\mathbf{y}^{4}\mathbf{x}^{-1}\mathbf{y}^{-1})$
$=\mathbf{x}^{12}.$